OpenStudy (anonymous):

convert this expression: 7^10-y = 4k into a logarithm

OpenStudy (amistre64):

7^(10-y) = 4k? log7(7^(10-y)) = log7(4k) (10-y) = log7(4) + log7(k) y= -[log7(4) + log7(k)] over (10) or are you looking for "k"?

OpenStudy (anonymous):

i want to change it into log

OpenStudy (amistre64):

log7(7) = 1 so the LHS becomes 1(10-y)

OpenStudy (anonymous):

\[b ^{y}=N\] \[= \log _{b}N=y\]where the base is b=7 the exponent is 10-y and N=4k so the answer is \[\log _{7}(4k)=10-y\] remember a log is an exponent

OpenStudy (anonymous):

thansk!